At a production level of 7,500 units, a project has earnings before interest and taxes of $48,310. Depreciation is $9,700, and fixed costs are $12,200. What is the variable cost per unit if the sales price per unit is $29.50?
Answer #1

EBIT = (Number of units) *(Selling price per unit - Variable cost) - Fixed cost - Depreciation

48310= 7500*(29.5 - Variable cost) - 12200 - 9700
48310= 7500*(29.5 - Variable cost) -21900
48310+21900= 7500*(29.5 - Variable cost)
70210= 7500*(29.5 - Variable cost)
70210/7500= 29.5 - Variable cost
9.361333333= 29.5 - Variable cost
Variable cost=29.5-9.361333333=20.13866667 or 20.14 (Rounded to two decimal places)